OpenStudy (anonymous):

If f(x) varies directly with x, and f(x) = 12 when x = 8, write the direct linear variation equation.

OpenStudy (akashdeepdeb):

***ALWAYS REMEMBER y = f(x)*** When they say that f(x) varies directly with x, it means that, As x increases, f(x) also increases, and as x decreases f(x) decreases. Take for example: f(x) = x  So in your example when they say that f(x) varies directly with x we say f(x) \(\alpha\) x Or f(x) = \(k.x\) [Where k is a constant of proportionality. So in your question it is given that when, f(8) = 12 or f(8) = k.8 [Because here, x = 8] Now if you find the value of 'k', then only can you write the equation, as y = k.x

OpenStudy (anonymous):

wow i do not get any of that

terenzreignz (terenzreignz):

Here's the deal... first, let's end confusion and let y = f(x) means replace f(x) with y. y looks a lot less menacing than f(x), don't you agree? We'll switch back later. Okay, ready?

OpenStudy (anonymous):

yes!

terenzreignz (terenzreignz):

When two values are in direct variation, it means their QUOTIENT is constant. So, since y and x are in direct variation, it means \[\Large \frac y x = k\] for some constant k, correct?

OpenStudy (anonymous):

okay..

terenzreignz (terenzreignz):

I don't like the uncertainty. If there's something you don't get, tell me :D

OpenStudy (anonymous):

no, no i get it so far

terenzreignz (terenzreignz):

Okay... what you must understand is that k is a CONSTANT. It doesn't change. All we have to do now is find out exactly what k is. Any idea how?

OpenStudy (anonymous):

no..

terenzreignz (terenzreignz):

Of course not (not meant as an insult ^_^) That's why you're here :D So, we know that \[\Large \frac yx = k\] and k is a constant that doesn't change... no matter what x and y are. It just so happens we were given specific values for x and y, do you know them? (mentioned in the question)

OpenStudy (anonymous):

y = 12, x = 8

terenzreignz (terenzreignz):

that's right. So we can plug those in, and we get (I wouldn't mind a drum roll) \[\Large \frac{12}{8}= k\] And suddenly, we have an exact value for k So, simplify it, what is k equal to?

OpenStudy (anonymous):

3/2?

terenzreignz (terenzreignz):

Good. So now, we know k = 3/2 And we know k doesn't change. So we have \[\Large \frac yx = \frac32\] Separate y. Make y stand alone on the left-side.

OpenStudy (anonymous):

y = 3/2x?

terenzreignz (terenzreignz):

\[\Large y = \frac32x\] That's right. Now you can replace y with f(x) and get your final answer ^_^

OpenStudy (anonymous):

wow thank you so much!

terenzreignz (terenzreignz):

Your final answer, post it, before thanking me XD

OpenStudy (anonymous):

f(x) = 3/2x (:

terenzreignz (terenzreignz):

And actually, @AkashdeepDeb has said pretty much the same things, only he said them all at once, which may have been a bit overwhelming ^_^ Kudos to him, nonetheless

OpenStudy (anonymous):

yeah thank you both, it was a bit confusing at first

terenzreignz (terenzreignz):

Now, remember these, these may be helpful in the future... If x and y are in DIRECT variation, it means their QUOTIENT is constant \[\Large \frac y x = k\] If x and y are in INVERSE variation, it means their PRODUCT is constant \[\large xy = k\] Get it? Got it? Good. :D

OpenStudy (anonymous):

yes thank you